Determine the [OH−] , pH, and pOH of a solution with a [H+] of 6.2×10−13M at 25 °C.


[OH−]= M



pH=



pOH=



Determine the [H+] , pH, and pOH of a solution with an [OH−] of 3.9×10 −13 M at 25 °C.


[H+]= M



pH=



pOH=



Determine the [H+] , [OH−] , and pOH of a solution with a pH of 2.28 at 25 °C.


[H+]= M



[OH−]= M



pOH=



Determine the [H+], [OH−] , and pH of a solution with a pOH of 4.08 at 25 °C.


[H+]= M



[OH−]= M



pH=




1
Expert's answer
2022-03-12T02:01:36-0500

pH + pOH = 14

pH = −log[H+]

pOH = −log[OH]

[H+] = 10-pH

[OH] = 10−pOH


(1) Determine the [OH] , pH, and pOH of a solution with a [H+] of 6.2×10−13 M at 25 °C.

Solution (1):

pH = −log[H+] = −log(6.2×10−13) = 12.21

pOH = 14 − pH = 14 − 12.21 = 1.79

[OH] = 10−pOH = 10−1.79 = 0.016 M = 1.6×10−2 M

Therefore,

[OH] = 1.6×10−2 M

pH = 12.21

pOH = 1.79


(2) Determine the [H+] , pH, and pOH of a solution with an [OH] of 3.9×10−13 M at 25 °C.

Solution (2):

pOH = −log[OH] = −log(3.9×10−13) = 12.41

pH = 14 − pOH = 14 − 12.41 = 1.59

[H+] = 10-pH = 10-1.59 = 0.0257 = 2.57×10−2 M

Therefore,

[H+] = 2.57×10−2 M

pH = 1.59

pOH = 12.41


(3) Determine the [H+] , [OH] , and pOH of a solution with a pH of 2.28 at 25 °C.

Solution (3):

[H+] = 10-pH = 10-2.28 = 5.25×10−3 M

pOH = 14 − pH = 14 − 2.28 = 11.72

[OH] = 10−pOH = 10−11.72 = 1.9×10−12 M

Therefore,

[H+] = 5.25×10−3 M

[OH] = 1.9×10−12 M

pOH = 11.72


(4) Determine the [H+], [OH] , and pH of a solution with a pOH of 4.08 at 25 °C.

Solution (4):

[OH] = 10−pOH = 10−4.08 = 8.3×10−5 M

pH = 14 − pOH = 14 − 4.08 = 9.92

[H+] = 10-pH = 10-9.92 = 1.2×10−10 M

Therefore,

[H+] = 1.2×10−10 M

[OH] = 8.3×10−5 M

pH = 9.92